x^4-5x^3+12x^2-5x-2
factorise it

Answer by uksraj(9)

f(x)=x^4-5x^3+12x^2-5x-2
f(-1)=-1^4-5(-1)^3+12(-1)^2-5(-1)-2
=1+5-12+5-1#0